# Forecasting and Manipulating the Forecasts of Others

### 18 Mar 2026 v2

@@ 3 Main results
Immediate from Proposition 3.6: the forward environment and backward adjoint are deterministic, so (3.9) is a deterministic variational problem. ∎
Remark 3.3 (Connection to rational inattention).
When $n=1$ the FOC for (3.9) reduces to the continuous-time analogue of Sims [27]; with $n>1$ the information wedge enters through $H^{X,i}$, so the marginal value of precision depends on the strategic environment—a more responsive opponent raises the return to filtering because better inference enables more effective belief manipulation.
@@
Under time-homogeneous primitives, response maps become time-invariant K(t,u)=k(tu)K(t,u)=k(t-u) and the fixed point becomes one in transfer functions, extending [12, 14, 24] to endogenous signals. Rational transfer functions correspond to finite-dimensional Markov sufficient statistics, so the pole structure of equilibrium kernels diagnoses whether an exact state-space reduction exists.
@@
$W=(W^{0},\dots,W^{n})$ primitive noise $E^{i},\;\Pi^{i}:=E^{i,\top}E^{i}$ selector / projector
@@
$(\bar{H}^{X,i},H^{X,i})$ physical adjoints $(\bar{H}^{k,i},H^{k,i})$ belief adjoints
$\mathcal{V}^{i}_{t}=\bar{\mathcal{V}}^{i}_{t}+\int_{0}^{t}\mathcal{V}^{i}_{t}(r)\,dW_{r}$ information wedge
4 Bilateral belief manipulation in a two-player game
To make the general theory concrete, we specialize to the simplest game that exhibits bilateral belief manipulation: two symmetric players with scalar state, opposing targets, and equal signal precision. This strips away all structure except the core strategic friction. Neither player observes the state directly; each receives a private diffusion signal of precision $p>0$. Because each player’s control enters the state dynamics, it reshapes the opponent’s signal and thereby the opponent’s posterior. The information wedge $\mathcal{V}^{i}_{t}$ of Theorem 3.2 prices this externality exactly; the mean wedge $\bar{\mathcal{V}}^{i}_{t}$ is the component we visualize below.
4.1 Primitives, objectives, and symmetry
Fix a horizon $[0,T]$. The state $X_{t}\in\mathbb{R}$ satisfies
$dX_{t}=\big(D_{t}^{1}+D_{t}^{2}\big)\,dt+dW_{t}^{0},\qquad X_{0}=x_{0},$ (4.1)
where $W^{0}$ is a standard Brownian motion. Player $i\in\{1,2\}$ does not observe $X$ directly; instead they observe the diffusion channel
with $(W^{0},W^{1},W^{2})$ mutually independent and $p>0$ representing root precision. Player $i$ chooses $D^{i}$ adapted to $\mathcal{F}_{t}^{i}=\sigma(Y_{s}^{i}:s\leq t)$.
There is no terminal cost in this worked example.
The model is invariant under the sign/label symmetry
$(X,\;Y^{1},\;Y^{2},\;D^{1},\;D^{2})\ \longmapsto\ (-X,\;Y^{2},\;Y^{1},\;-D^{2},\;-D^{1}).$
[…]
Figure 2: Information-to-action feedback in the worked example. Actions move $X_{t}$ and shift opponents’ beliefs, so equilibrium mean actions depend on precision $p$.
4.2 Numerical equilibrium and kernel visualization
We now present the numerical equilibrium obtained from the Picard iteration described above. All figures below are generated from the converged kernel profile $D^{1}$, its primitive-noise representation $\mathcal{D}^{1}$, the state kernel $X$, and the filtering kernel $F^{1}$.
Computation.
Equilibrium is computed by iterating the best-response operator on the deterministic impulse-response system, discretized on a uniform grid of $N=40$ points; no Monte Carlo simulation is needed. At each iterate the forward environment $(X,\widetilde{X}^{i},F^{i})$ is recomputed from the closure equations, the backward adjoint $H^{X,i}$ is solved by explicit backward Euler, and the updated policy is blended into the current profile. Parameters: $T=1$, $x_{0}=0$, $r=0.1$, $G^{XX,i}(T)=0$. The outer residual reaches $10^{-5}$ in $19$ iterations( ${\approx}\,10$ ms in C++). At longer horizons the binding constraint is representational (the filtering kernel becomes more singular) rather than economic.
Convergence of the fixed point.
The outer Picard residual decays geometrically, confirming convergence of the fixed-point scheme in this symmetric benchmark; we use the iteration as a constructive computation rather than an existence proof.
Impulse-response viewpoint.
The key distinction is between $D^{1}_{t}(s)$ (response to a shock as inferred through the filter).
(a) Primitive-noise state kernel $X_{t}(s)$.
(b) Equilibrium noise-state feedback kernel $D^{1}_{t}(s)$.
Figure 3: Equilibrium kernel time-slices, shown channel-by-channel for $(W^{0},W^{1},W^{2})$ with curves colored by evaluation time $t$. (a) Both players jointly stabilize older fundamental shocks (decay in the $W^{0}$ panel); observation-noise channels show smaller, opposed effects mediated by the belief loop. (b) The system-noise channel dominates; observation-noise channels exhibit curvature reflecting drift-based inference.
4.3 Mean actions, separation failure, and bilateral belief manipulation
Since the stochastic integral in $D_{t}^{i}=\bar{D}^{i}_{t}+\int_{0}^{t}D^{i}_{t}(u)\,d_{u}\widehat{W}_{t}^{i}(u)$ has zero unconditional mean, $\mathbb{E}[D_{t}^{i}]=\bar{D}^{i}_{t}$ isolates the deterministic tug-of-war over the target $\pm 1$, while the kernel term captures impulse responses to posterior shocks. Under perfect information, the mean policy does not depend on signal precision—a certainty-equivalence property. Decentralized information breaks this: changing precision changes the speed at which opponents update, giving each player a reason to act on posteriors rather than just on the physical state.
The information wedge as a dynamic rent density.
In the general theory (Theorem 3.2), the backward equation for $\bar{H}^{X}$ contains opponent-filter coupling terms that vanish under perfect information or in the single-player case. Here these terms aggregate into
$\bar{\mathcal{V}}^{1}_{t}=P^{2}(t)\int_{0}^{t}\widetilde{X}^{2}_{t}(z)\cdot\bar{H}^{2}_{t}(z)\,dz,$
which enters the mean adjoint equation as $\tfrac{d}{dt}\bar{H}^{X}_{t}=-(\bar{X}_{t}-1)-\bar{\mathcal{V}}^{1}_{t}$. The unresolved kernel $\widetilde{X}^{2}_{t}(z)$ measures how much of the $z$-th shock remains unresolved by player 2 at time $t$; the belief-adjoint $\bar{H}^{2}_{t}(z)$ measures the shadow value to player 1 of that shift. Their inner product $\bar{\mathcal{V}}^{1}_{t}$ prices the mean channel of bilateral belief manipulation at time $t$; Figures 56 visualize this object across precision asymmetries.
Figure 7: Equilibrium costs under private signals (dashed) and pooled signals (solid) with $p_{1}=3$ fixed and $p_{2}$ varying. Left: competitive targets ($\theta_{1}=1$, $\theta_{2}=-1$). Right: cooperative targets ($\theta_{1}=\theta_{2}=0$). Pooling is Pareto-improving in both cases, but the gain is an order of magnitude larger under competition, where it eliminates the bilateral belief-manipulation externality priced by $\mathcal{V}^{i}_{t}$. Blue: player 1; red: player 2.
Information pooling as a Pigouvian intervention.
Figure 7 compares equilibrium costs under private signals with those under a pooled signal of precision $P^{1}+P^{2}$, for both competitive ($\theta_{1}=1$, $\theta_{2}=-1$) and cooperative ($\theta_{1}=\theta_{2}=0$) targets. Pooling is Pareto-improving in both cases, but the gains differ by an order of magnitude: roughly 0.020.020.050.05 under cooperation versus 0.30.30.50.5 under competition. Since the estimation benefit of pooling enters identically regardless of the target structure, the difference isolates the strategic externality channel priced by $\mathcal{V}^{i}_{t}$: nearly all of the welfare gain from mandatory disclosure comes from eliminating bilateral belief manipulation, not from improving state estimation.
5 Market microstructure as a special case
The two-player example operates within the quadratic-cost framework of Section 2. We now show that the noise-state technique embeds the Kyle–Back model with no additional machinery, despite traders maximizing expected profits against the market maker’s price rather than minimizing a quadratic objective. The embedding works because $\lambda(t)>0$ provides the strict convexity needed for the best-response closure, the filtering closure applies verbatim to the stacked observation channel, and the backward system of Appendix D yields the optimality condition. We derive the equilibrium characterization for a single informed trader and identify the two channels through which belief manipulation operates; the extension to multiple asymmetrically informed traders follows the same backward system and is a natural application of the general theory.
@@ 5.1 Model and embedding
Adjoin a fundamental-value Brownian motion $W^{V}$ independent of $(W^{0},\dots,W^{n})$, with selector $E^{V}$ and projector $\Pi^{V}$. The fundamental is $V_{t}:=\sigma_{V}W_{t}^{V}$. All players observe aggregate order flow $dZ_{t}=\sum_{j}D_{t}^{j}\,dt+\sigma_{0}\,dW_{t}^{0}$; trader $i$ additionally observes a private signal $dS_{t}^{i}=\sqrt{p_{i}}\,V_{t}\,dt+dW_{t}^{i}$. The market maker observes only $Z$ and posts the competitive price $P_{t}:=\mathbb{E}[V_{T}\mid\mathcal{F}_{t}^{0}]=\sigma_{V}\int_{0}^{t}E^{V}d_{u}\widehat{W}_{t}^{0}(u)$. Trader $i$ maximizes $\pi^{i}:=\mathbb{E}[\int_{0}^{T}D_{t}^{i}(V_{T}-P_{t})\,dt]$. By symmetry $W\mapsto-W$, all mean terms vanish.
Embedding.
Rescaling the order-flow channel to unit noise, trader $i$’s stacked observation is
$d\!\begin{pmatrix}Z_{t}/\sigma_{0}\\ S_{t}^{i}\end{pmatrix}=\underbrace{\begin{pmatrix}\sigma_{0}^{-1}\\[2.0pt] \sqrt{p_{i}}\end{pmatrix}}_{\sqrt{P^{i}}}\!\!\begin{pmatrix}\text{demand drift}\\ \sigma_{V}E^{V}W_{t}\end{pmatrix}\!dt\;+\;\underbrace{\begin{pmatrix}E^{0}\\ E^{i}\end{pmatrix}}_{\mathcal{E}^{i}}\!dW_{t},$ (5.1)
which is the observation equation of Section 2 with $\sqrt{P^{i}}=(\sigma_{0}^{-1},\sqrt{p_{i}})^{\top}$ and $E^{i}\to\mathcal{E}^{i}$; the market maker has $\sqrt{P^{0}}=\sigma_{0}^{-1}$, $\mathcal{E}^{0}=E^{0}$.
Kyle’s lambda.
@@
$dP_{t}=\lambda(t)\,dI_{t}^{0},\qquad\lambda(t):=\sigma_{V}\!\int_{0}^{t}E^{V}\widetilde{C}^{0}(t,u)^{\top}du,$ (5.2)
where $\widetilde{C}^{0}$ is the market maker’s unresolved drift kernel under the projection decomposition (A.12).
Remark 5.1 (Quadratic structure of the profit).
Fix opponents’ strategy maps. Since the market maker’s filtering gains are deterministic, the price decomposes as $P_{t}=P_{t}^{-i}+\int_{0}^{t}\ell(t,s)\,D_{s}^{i}\,ds$, where $P_{t}^{-i}$ is a linear process independent of $D^{i}$ and $\ell(t,s)$ is a deterministic kernel. Symmetrically, the future execution cost of a unit impulse at time $t$ is $\int_{t}^{T}D_{\tau}^{i}\,\delta P_{\tau}\,d\tau$, where $\delta P_{\tau}$ propagates forward through the deviation system with deterministic coefficients. The profit $\pi^{i}=\mathbb{E}[\int_{0}^{T}D_{t}^{i}(V_{T}-P_{t})\,dt]$ is therefore concave quadratic in $D^{i}$, when $\lambda(t)>0$ provides strict concavity.
5.2 Equilibrium characterization
The optimality condition from a unit demand impulse at $t$ equates the conditional mispricing to the integrated impact cost (Appendix D):
$\underbrace{\sigma_{V}E^{V}\,\partial_{u}\bigl(\widehat{W}_{t}^{i}-\widehat{W}_{t}^{0}\bigr)(u)}_{H^{i,\mathrm{val}}_{t}(u):\;\text{information rent density}}\;=\;\underbrace{\int_{t}^{T}\delta P_{\tau}\,D^{i}_{\tau}(u)\,d\tau}_{H^{i,\mathrm{imp}}_{t}(u):\;\text{impact cost density}},$ (5.3)
where $\delta P_{\tau}$ is the price perturbation for an impulse at $t$. Differentiating both sides in $t$ and solving for $D^{i}_{t}(u)$ yields a closed backward ODE:
$\lambda(t)\,D^{i}_{t}(u)=-\frac{d}{dt}H^{i,\mathrm{val}}_{t}(u)+\int_{t}^{T}\partial_{t}\delta P_{\tau}\,D^{i}_{\tau}(u)\,d\tau,$ (5.4)
with terminal condition $D^{i}_{T}(u)=\lambda(T)^{-1}H^{i,\mathrm{val}}_{T}(u)$.
@@ 5.2 Equilibrium characterization
This requires that aggregate order flow remains informative about the fundamental at every date, the economically relevant case since temporary uninformativeness would present an arbitrage opportunity. The condition is self-reinforcing: positive $\lambda$ induces finite trading rates, which maintain the drift informativeness that sustains $\lambda>0$. It plays the role of $G^{DD,i}\succ 0$ (Lemma 3.4), ensuring strict convexity of the best-response problem on the domain of the fixed-point operator.
The derivative $\frac{d}{dt}H^{i,\mathrm{val}}$ and the deviation system determining $\delta P_{\tau}$ are given in Appendix D.
5.3 Two channels of belief manipulation
Decompose $\delta P_{\tau}$ into $\delta P_{\tau}^{\mathrm{dir}}$ (computed with $D^{k}\equiv 0$, i.e. the market maker’s response alone) and $\delta P_{\tau}^{\mathrm{ind}}:=\delta P_{\tau}-\delta P_{\tau}^{\mathrm{dir}}$ (opponent-mediated). Since $\delta P_{\tau}^{\mathrm{dir}}=\lambda(t)$ for all $\tau\geq t$, the impact adjoint splits into:
  1. (i)
    $\mathcal{V}^{i,\mathrm{price}}(t,u)=\lambda(t)\int_{t}^{T}D^{i}_{\tau}(u)\,d\tau$: instantaneous impact times the integrated future policy kernel. This is the standard Kyle channel—the trader internalizes how each demand impulse moves the price.
  2. (ii)
    $\mathcal{V}^{i,\mathrm{opp}}(t,u)=\int_{t}^{T}\delta P_{\tau}^{\mathrm{ind}}\,D^{i}_{\tau}(u)\,d\tau$: opponent-mediated impact, arising because opponents also observe order flow and adjust their trading, which feeds back into prices.
Back’s model.
When $n=1$ and $p_{1}=0$ (Back’s model), the opponent channel vanishes ($\delta P_{\tau}^{\mathrm{ind}}\equiv 0$) and (5.3) reduces to Back’s integral equation for the optimal trading rate. With $n\geq 2$ and $p_{i}>0$ both channels are active: each trader tempers demand not only to limit direct price impact but to prevent opponents from inferring and front-running through shared order flow. Guess-and-verify cannot close this system; the noise-state technique handles it with no additional machinery.
@@ 6 Conclusion
Conditioning on primitive Brownian shocks collapses the Townsend belief hierarchy onto deterministic impulse-response maps, yielding the first exact equilibrium characterization of finite-player continuous-time LQG games with endogenous signals. The information wedge $\mathcal{V}^{i}_{t}$ prices bilateral belief manipulation and vanishes when signals are exogenous. The two-player example shows that nearly all welfare gains from disclosure come from eliminating this channel; the Kyle–Back embedding shows the technique extends beyond quadratic costs, identifying a second channel through opponent inference via shared order flow that applies to asymmetric precisions and continuous private learning.
@@
The main open question is whether the deterministic impulse-response class is without loss of generality. A natural program proceeds in three steps: (i) a Blackwell reduction showing that under quadratic costs the best response depends on the posterior $\pi_{t}^{i}:=\mathcal{L}(W_{[0,t]}\mid\mathcal{F}_{t}^{i})$ only through its mean and covariance; (ii) a belief-of-belief collapse in which player $i$’s estimate of $\pi_{t}^{k}$ factors through a deterministic operator—the posterior analogue of the kernel maps $(F^{i},\widetilde{X}^{i})$; (iii) a contraction argument on the variance functional $V(\mathbf{D}):=\sum_{i}\mathbb{E}\int\!\!\int\|D^{i}(t,s,\pi_{t}^{i})-\mathbb{E}[D^{i}(t,s,\pi_{t}^{i})]\|^{2}\,ds\,dt$, which vanishes exactly in the deterministic impulse-response class and which iterated best responses are expected to contract. Beyond this, stochastic coefficients (Remark C.2), graph-structured interactions, and global uniqueness via the dissipative structure of $F^{i}$ are natural extensions.

### 23 Mar 2026 v3

@@ 4 Main results
With $n\geq 2$ the information wedge couples the costate to belief adjoints $H^{k,i}_{t}(u)$, which depend on $\widetilde{X}^{k}_{t}(u)$—the portion of the state response opponent $k$ has not yet learned. This introduces $(t,u)$-dependence that cannot factor through $X_{t}(u)$. Perturbing $P^{i}$ changes $\widetilde{X}^{k}$ through the equilibrium, changing $H^{k,i}$ and hence the policy kernel independently of the state estimate. The factorization holds if and only if $\mathcal{V}^{i}_{t}\equiv 0$. ∎
Remark 4.5 (Risk-sensitive extension).
The framework extends to the entropic risk measure $J^{i}_{\theta_{i}}:=\theta_{i}^{-1}\log\mathbb{E}[e^{\,\theta_{i}\mathcal{C}^{i}}]$: the best-response FOC is identical to Corollary 4.5 with $\widehat{W}_{t}^{i}$ replaced by a risk-adjusted noise-state that tilts conditional means and rescales conditional precision through the quadratic cost kernel (Appendix B.6, Theorem B.1).
@@
Under time-homogeneous primitives, response maps become time-invariant K(t,u)=k(tu)K(t,u)=k(t-u) and the fixed point becomes one in transfer functions, extending [14, 16, 27] to endogenous signals, where the transfer function of the equilibrium filter is itself an equilibrium object. Rational transfer functions correspond to finite-dimensional Markov sufficient statistics, so the pole structure of equilibrium kernels diagnoses whether an exact state-space reduction exists.
@@
$W=(W^{0},\dots,W^{n})$ primitive noise $E^{i},\;\Pi^{i}$ selector / projector
@@
$H^{X,i}_{t}$ physical costate (linear process) $H^{k,i}_{t}(u)$ belief adjoints (linear process)
$\mathcal{V}^{i}_{t}$ information wedge (linear process)
5 Trading under asymmetric informedness
The noise-state technique extends to strategic trading with asymmetric private learning, a setting where guess-and-verify breaks down once traders differ in precision or learn continuously from private sources. The price each trader faces decomposes into an exogenous component and a history-linear function of their own demand, with Kyle’s lambda playing the role of $G^{DD,i}$; the quadratic structure is distributed across time, so the first-order condition yields an integral equation whose time derivative pins down the optimal trading rate.
@@ 5.1 Model and embedding
There are $n$ informed traders and a competitive market maker. A fundamental $V_{t}:=\sigma_{V}W_{t}^{V}$ is driven by a Brownian motion $W^{V}$ independent of $(W^{0},\dots,W^{n})$, with selector $E^{V}$. All players observe aggregate order flow $dZ_{t}=\sum_{j}D_{t}^{j}\,dt+\sigma_{0}\,dW_{t}^{0}$; trader $i$ additionally observes a private signal $dY_{t}^{i}=\Gamma^{i}(t)\,V_{t}\,dt+dW_{t}^{i}$. The market maker observes only $Z$ and posts the competitive price $P_{t}:=\mathbb{E}[V_{T}\mid\mathcal{F}_{t}^{0}]$. Trader $i$ maximizes expected profit $\pi^{i}:=\mathbb{E}[\int_{0}^{T}D_{t}^{i}(V_{T}-P_{t})\,dt]$. By the symmetry $W\mapsto-W$, all mean terms vanish.
Stacking trader $i$’s two observation channels into a single equation recovers the observation model of Section 3 (Appendix E), so the filtering machinery of Appendix A applies verbatim.
Price decomposition.
Fix opponents’ noise-state linear strategy maps. The market maker’s filtering gains are deterministic, so the price decomposes as
$P_{t}\;=\;P_{t}^{-i}\;+\;\int_{0}^{t}\ell(t,s)\,D_{s}^{i}\,ds,$ (5.1)
where $P_{t}^{-i}$ is a linear process independent of $D^{i}$ off equilibrium (the price that would obtain if trader $i$ submitted zero demand) and $\ell(t,s)$ is a deterministic kernel encoding how each unit of past demand at time $s$ has moved the current price by time $t$. The profit becomes
$\pi^{i}=\mathbb{E}\!\int_{0}^{T}D_{t}^{i}\underbrace{(V_{T}-P_{t}^{-i})}_{\text{exogenous mispricing}}\,dt\;-\;\mathbb{E}\!\int_{0}^{T}D_{t}^{i}\underbrace{\int_{0}^{t}\ell(t,s)\,D_{s}^{i}\,ds}_{\text{endogenous price impact}}\,dt.$
The first term is linear in $D^{i}$; the second is a quadratic form in the history of $D^{i}$ through the deterministic kernel $\ell$: $(V_{T}-P_{t}^{-i})$ plays the role of $G^{X,i}$, and the integrated price-impact term plays the role of $D^{i,\top}G^{DD,i}D^{i}$, distributed across time rather than concentrated at each instant.
Kyle’s lambda and strict concavity.
@@
$dP_{t}=\lambda(t)\,dI_{t}^{0}.$ (5.2)
@@
This is the economically relevant case: if aggregate order flow were temporarily uninformative ($\lambda=0$), traders could trade at zero price impact—an arbitrage opportunity that would restore informativeness. The condition plays the role of $G^{DD,i}(t)\succ 0$ in the LQG framework (Lemma 4.4), ensuring strict concavity of the profit in $D^{i}$ and hence necessity and sufficiency of the first-order condition, just through the temporally distributed quadratic form rather than pointwise at each $t$.
5.2 Best-response characterization
Fix player $i$ and suppose every opponent $k\neq i$ uses a noise-state linear strategy as in Section 4. Because the quadratic structure is cross-temporal, the first-order condition does not directly yield $D^{i}_{t}(u)$; instead, the optimality condition equates the conditional mispricing to the integrated future impact cost (Appendix E):
$\underbrace{\,\widehat{V}^{i}_{t}-P_{t}}_{H^{i,\mathrm{val}}_{t}(u):\;\text{information rent density}}\;=\;\underbrace{\mathbb{E}\left[\int_{t}^{T}l(\tau,t)\,D^{i}_{\tau}(u)\,d\tau|\mathcal{F}^{i}_{t}\right]}_{H^{i,\mathrm{imp}}_{t}:\;\text{impact cost density}},$ (5.3)
where $\delta P_{\tau}$ is the price perturbation propagated by the deviation system (E.1)–(E.3). The left side is trader $i$’s informational advantage over the market maker; the right side is the present value of price distortion from a marginal unit of demand. Since the integral runs over $[t,T]$, it does not isolate $D^{i}_{t}(u)$—only the entire future trading plan weighted by price impact.
Differentiating both sides in $t$, the Leibniz boundary term produces $\lambda(t)D^{i}_{t}(u)$—where $\lambda(t)$ enters as the endogenous control cost and $D^{i}_{t}(u)$ first appears explicitly. Solving:
$\lambda(t)\,D^{i}_{t}=-\frac{d}{dt}H^{i,\mathrm{val}}_{t}+\mathbb{E}\left[\int_{t}^{T}\frac{\partial l}{\partial t}(\tau,t)\,D^{i}_{\tau}\,d\tau|\mathcal{F}^{i}_{t}\right],$ (5.4)
with terminal condition $D^{i}_{T}(u)=\lambda(T)^{-1}H^{i,\mathrm{val}}_{T}(u)$. The derivative $\frac{d}{dt}H^{i,\mathrm{val}}$ and the deviation system determining $\delta P_{\tau}$ are given in Appendix E.
@@ 6 Conclusion
Conditioning on primitive Brownian shocks collapses the Townsend belief hierarchy onto deterministic impulse-response maps, yielding the first exact equilibrium characterization of finite-player continuous-time LQG games with endogenous signals. The information wedge $\mathcal{V}^{i}_{t}$ prices bilateral belief manipulation and vanishes when signals are exogenous. The two-player example shows that nearly all welfare gains from disclosure come from eliminating this channel; the Kyle–Back embedding shows the technique extends beyond quadratic costs to strategic trading with endogenous price impact, asymmetric precisions, and continuous private learning.
@@
The main open question is whether the deterministic impulse-response class is without loss of generality. A natural program proceeds in three steps: (i) a Blackwell reduction showing that under quadratic costs the best response depends on the posterior $\pi_{t}^{i}:=\mathcal{L}(W_{[0,t]}\mid\mathcal{F}_{t}^{i})$ only through its mean and covariance; (ii) a belief-of-belief collapse in which player $i$’s estimate of $\pi_{t}^{k}$ factors through a deterministic operator—the posterior analogue of the kernel maps $(F^{i},\widetilde{X}^{i})$; (iii) a contraction argument on the variance functional $V(\mathbf{D}):=\sum_{i}\mathbb{E}\int\!\!\int\|D^{i}(t,s,\pi_{t}^{i})-\mathbb{E}[D^{i}(t,s,\pi_{t}^{i})]\|^{2}\,ds\,dt$, which vanishes exactly in the deterministic impulse-response class. The precision optimization (Appendix C.2) reveals that multi-player rational inattention couples the filtering and control problems that separate in single-agent settings: the marginal value of precision depends on the equilibrium through the state kernel. Extending the framework to rational inattention requires care: varying the gain $\Gamma^{i}$ directly changes the filtration in a non-monotone way, breaking the envelope argument. The difficulty disappears when the control variable is the portfolio of independent sources monitored (a Brownian sheet formulation), since each additional source strictly enlarges the filtration. The total precision decomposes as $P^{i}=P^{i}_{0}+\tau^{i}$, the marginal value of a source is the value of optimally exploiting its independent increment, and the envelope theorem applies cleanly (Appendix C). Overlapping source portfolios endogenize cross-player noise correlations, amplifying the belief-manipulation channel priced by $\mathcal{V}^{i}_{t}$.
An alternative to the mean-field limit is games on vertex-transitive graphs, where graph symmetry permits a representative-agent reduction while preserving the information wedge. The natural conjecture is that equilibrium impulse-response maps inherit the graph symmetry, reducing the fixed point to a single node coupled to its neighborhood—interpolating between the mean-field limit (complete graph, vanishing edge weights, $\bar{\mathcal{V}}^{i}_{t}=0$) and the finite-$n$ case. Stochastic coefficients (Remark D.2) and global uniqueness are natural next steps; the dissipative structure of Lemma D.3 suggests sufficient regularity to preclude multiplicity, but a proof requires monotonicity estimates beyond the Grönwall bounds used here.